Whats (5x)^(-1/2)=10? Steps please
Answer by lwsshak3(11628)

Whats (5x)^(-1/2)=10
(5x)^(-1/2)=10
1/(5x)^(1/2)=10
√(5x)=1/10
square both sides
5x=1/100
x=1/500